given the function f(x)=2x+10 find each of the following
f(9),f(-7),f(0)

Answer by Maths68(1474)

f(9)
f(x)=2x+10
f(9)=2(9)+10
f(9)=18+10
f(9)=28


f(-7)
f(x)=2x+10
f(-7)=2(-7)+10
f(-7)=-14+10
f(-7)=-4


f(0)
f(x)=2x+10
f(0)=2(0)+10
f(0)=0+10
f(0)=10